OpenStudy (anonymous):

f(x) = 4x-1/2 then f^-1(x)=????

OpenStudy (anonymous):

interchange y and x and solve for y

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

wats the 2nd equation?

OpenStudy (anonymous):

okay rewrite as y = 4x - 1/2. then switch y and x x = 4y - 1/2. then solve for y

OpenStudy (anonymous):

but there is no y in the problem

OpenStudy (anonymous):

replace f(x) with y

OpenStudy (anonymous):

ohhhhhh!!! ok:D makes sense now

OpenStudy (anonymous):

so y = 4y-1/2?

OpenStudy (anonymous):

wait i am sorry do you mind solving it for me and showing it please....

OpenStudy (anonymous):

y = 4x - 1/2 switch x = 4y - 1/2 solve for y : x + 1/2 = 4y (1/4)x + 1/8 = y so f^-1(x) = (1/4)x +1/8

OpenStudy (anonymous):

thank you! =)